# Stability and dynamics of massive vortices in two-component Bose-Einstein condensates

### v1

@@ III.2 Vortex-bright dynamics
We now leverage the ODE reduction dynamics presented in Sec. II.2 to describe the dynamics of the VB in the original PDE model. To generate rotational dynamics of the VB state at the level of the original GP model (1), we displace the stationary solutions away from the origin and seed the waveform into a ground state cloud according to the following procedure. First we compute, using our fixed-point algorithm, the ground state cloud $\psi_{1}=\psi_{\rm gs}$ for a given chemical potential value $\mu_{d}$ and a fixed value of the inter-species coupling $g_{12}$ (while we keep the trap strength constant $\Omega=0.1$ throughout this analysis). For convenience we choose a starting point with $x(0)=x_{0}$ and $y(0)=0$. Next, we interpolate the local chemical potential $\mu_{x_{0}}=|\psi_{\rm gs}(x_{0},0)|^{2}$ where the VB will be seeded. The VB solution with $\mu_{d}=\mu_{x_{0}}$ for a chosen $\mu_{b}$ value is prepared and subsequently imprinted at the location $x_{0}$ onto the ground state cloud $\psi_{\rm gs}$. The resulting displaced structure is propagated according to the dynamical equation (1). Note that to avoid unnecessary oscillations of the background cloud induced by the imprinted VB, the initial condition is “kicked” with a suitable (small) velocity $k$ by boosting $\psi_{2}$ with $e^{iky}$ in order to eliminate any linear momentum for the dark cloud. Our study of the resulting VB trajectories will focus mostly on parameter regimes where flowering trajectories are obtained.
@@
For the relatively small $g_{12}$ values shown in Fig. 1, the resulting PDE trajectories after displacement away from the origin are observed to be approximately circular with additional wobbling motion that does not genuinely resemble flowering trajectories. Nonetheless, for the relatively larger $g_{12}$ values shown in Fig. 2 and for large enough $\mu_{b}$ values, the PDE trajectories result in flowering motion. It is in this regime that we will use the ODE reduction provided in Sec. II to match the full GP VB dynamics. For example, keeping $\Omega=0.1$ and for $g_{12}=1.6$, $\mu_{d}=4.25$, and $x_{0}=2.5$, GP flowering trajectories are seen for $\mu_{b}\gtrsim 4.4$ while spiral-outs are seen for $\mu_{b}<4.4$. This is consonant with the stability results depicted in Fig. 5. However, it is important to note that the stability results of Fig. 5 correspond to a VB that is centered at the origin while we are now seeding VB structures away from the center. For instance, we note that within the $\mu_{b}$-interval $4.44.383<\mu_{b}<$ 4.4 stationary (i.e., centered at the origin) solutions are stable, but the displaced state starting at $x_{0}=2.5$ results in a spiral-out motion. This evidences that this displacement is relatively far away from the origin for the stability analysis at the origin to be valid. Nonetheless, we have confirmed that, in the unstable region $\mu_{b}\lesssim 4.383$, slightly perturbed solutions from the origin tend to oscillate with frequencies given by Im$(\nu)$.
@@
To compare the ODE reduction with the full GP dynamics in the stable regime of Fig. 5 we also depicted therein the value for $\dot{\theta}$ from the ODE model overlaid to demonstrate the frequency matching between the PDE and ODE models. This matching is done according to the following procedure. For a given flowering trajectory in the PDE model within appropriate parameter regimes, to obtain the corresponding ODE parameters we first identify those eigenvalues where the Krein signature bifurcates from zero to positive and negative values. The eigenvalues in Fig. 5 are color coded using (blue, black, red) points when the Krein sign is (negative, zero, positive). In particular, for a given $\mu_{b}$ value, we use the upper branch of the bifurcated eigenvalues to obtain the value $\dot{\theta}={\rm Im}(\nu)$. Then, by Eq. (18) we write the corresponding ODE $\Lambda$ parameter as:
$\Lambda=2[\rho(r_{0})\dot{\theta}-M_{0}\dot{\theta}^{2}].$ (26)
@@
Figure 9: VB trajectories for $g_{12}=1.6$, $M=0.01$, $\mu_{d}=4.25$, and $\mu_{b}=4.56$. The VB is seeded at $x_{0}=2.5$ corresponding to a local chemical potential $\mu_{x_{0}}\approx 4.22$. To match the ODE trajectories we use $M_{0}\approx 18.9$. The top-left panel shows the resulting trajectory with initial velocities $k_{y}=0$ in the PDE (blue curve) and $\dot{y}(0)\approx 0.01$ in the ODE (red curve) with a resulting angular momentum of $\ell\approx 8.9$. For this value of $\ell$, the effective potential $V_{\rm eff}(r)$ in Eq. (16) is plotted in the bottom-right panel (blue). Three black dots are plotted on $V_{\rm eff}(r)$ corresponding to its minimum and the minimum and maximum radii of the trajectory. From these values we have an effective flower energy of $\Delta E\approx 0.002$ according to Eq. (28). Parameters are the same in the top-right panel except with initial velocities $k_{y}=0.1$ in the PDE and $\dot{y}(0)\approx 0.06$ in the ODE, with $\ell\approx 8.92$ giving the effective potential in the bottom-right panel (red) with $\Delta E\approx 0.01$. Similarly, the bottom-left panel corresponds to $k_{y}=0.15$ in the PDE and $\dot{y}(0)\approx 0.08$ in the ODE, while $\ell\approx 8.93$ giving the effective potential in the bottom-right panel (yellow) with $\Delta E\approx 0.03$. The thin black circles in the trajectory plots correspond to the circular orbits at the minima of the effective potentials.
@@
Figure 10: Similar to Fig. 9 with the same parameters except here $\mu_{b}\approx 4.47$, $M=0.02$, and $M_{0}\approx 37.0$. The top left panel has initial velocities $k_{y}=0$ in the PDE, and $\dot{y}(0)\approx 0.01$ in the ODE, $\ell\approx 8.9$, and $\Delta E\approx 0.006$. For this $\ell$, $V_{\rm eff}(r)$ is in the bottom right panel (blue). For the top right panel $k_{y}=0.1$ in the PDE, $\dot{y}(0)\approx 0.06$ in the ODE, with $\ell\approx 8.95$ to give the effective potential in the bottom right panel (red) with $\Delta E\approx 0.016$. Similarly in the bottom left panel $k_{y}=0.15$ in the PDE, $\dot{y}(0)\approx 0.08$ in the ODE, $\ell\approx 8.97$ and the corresponding effective potential is in the bottom right (yellow) with $\Delta E\approx 0.056$. Once again the thin black circles in the trajectory plots represent the circular orbits at the minima of the effective potentials.
@@
Figure 8 depicts the corresponding ODE parameters across a range of $\mu_{b}$ values to the right of the Hamiltonian Hopf bifurcation point. The blue dashed line $P_{2}/\pi$ depicts the estimate for our ODE parameter $M_{0}$ according to Ref. Ruban (2022) [note their $(\tilde{M},M_{br})$ corresponds to our $(M_{0},P_{2})$ respectively as their Eq. (29) agrees with our Eq. (10); see also their Eq. (24) giving $M_{0}=M_{br}/\pi=P_{2}/\pi$]. The match of the relevant curves in Figure 8 is very good, revealing only a slight relative discrepancy of no more than $3.5\%$ for the estimate of $M_{0}$. On the other hand, the red dashed line corresponds to the precessional frequency of a vortex in the absence of the bright component given by $\Lambda=2\omega_{\rm pr}^{0}\mu_{d}$ obtained from Ref. Middelkamp et al. (2010) as follows. Taking our Eq. (10) with $M_{0}=0$ near the origin we obtain
@@
Comparing this reduced system to Eqs. (3) and (4) of Ref. Middelkamp et al. (2010) gives $\Lambda=2w_{\rm pr}^{0}\mu_{d}\equiv\Omega^{2}\ln(8.88\mu_{d}/\Omega)$. The relatively large discrepancy between $\Lambda$ and the theoretical estimate may stem from the fact that the latter, rather than being a true estimate for the precessional frequency of the (massive) VB, it is just the precessional frequency corresponding to a vortex in the absence of the bright component. Note that, for this region of parameters (see Fig. 2), as $\mu_{b}$ increases the mass of the bright component decreases and thus, correspondingly, the $\Lambda$ curve progressively approaches the (red dashed) $\Lambda=2\omega_{pr}^{0}\mu_{d}$ non-filled vortex benchmark.
@@ IV Summary and outlook
In the present work we have explored the topic of the dynamics of VB solitary waves deep within the immiscible regime. Our motivation in revisiting this subject stemmed from the renewed experimental ability to monitor such systems which creates the potential for connecting a particle-based understanding with a detailed analysis of the full PDE model. In that vein, we explored equilibrium solutions in this deeply immiscible regime (while some of the earlier studies of the present authors had confined themselves to regimes close to, e.g., hyperfine states of 87Rb where the $g_{ij}$ coefficients are nearly equal Law et al. (2010)). In probing the excitations around the VB equilibrium state, we found the possibility of two frequencies of precession. One of these represents the frequency of precession that is well-familiar from a single-component vortex in a trap. The second frequency can be associated with the oscillation of the VB structure inside an effective radial potential (along the radial direction). The combination of these two frequencies gives rise to the potential for epicyclic motion, which, under suitable conditions discussed herein, can lead to “flowering”, i.e., to flower-like curves of the dynamics of the VB pattern. This intriguing feature, while observed and discussed previously Richaud et al. (2021); Ruban (2022) is here intimately connected between the ODE and PDE dynamics (both at the level of linearization and at that of fully nonlinear dynamical evolution). Our analysis enables both a reasonable approximate analytical description of the relevant motion and key byproducts thereof, such as the critical oscillation amplitude leading to flowering as a (simple) function of the precessional and radial oscillation frequencies. Among the additional intriguing features identified herein, we highlight the (oscillatory) instability of the VB structure when the relevant two frequencies collide. The result of the corresponding dynamics was found to be particularly unusual, in our experience of such systems. Indeed, as the VB embarks on a growing spiral oscillation, it becomes a patch of vorticity that gets elongated and eventually breaks into smaller patches carrying multiple localized phase jumps inside an extended region of vanishing density (in the so-called dark component) in a way reminiscent of persistent currents Wright et al. (2013).
@@
While we considered a number of diagnostics (such as angular momentum-based ones) and visualizations (such as the 3D isocontours of vorticity) of this phenomenon, clearly it merits further study. An additional natural theme of study is to incorporate the (two or more) VB interactions to the model systematically developed herein, so as to examine the potential of formation of lattices (crystals) of such patterns. Understanding the modes of such crystals, their potential instabilities and dynamics promises to be an interesting avenue for the future. Furthermore, the states of interest can be generalized to three spatial dimensions where the natural analogue is the vortex-ring-bright (filled core) solitary wave Ruban et al. (2022); of course, vortex-line-bright structures exist too Charalampidis et al. (2016). A systematic particle-based understanding of such 3D settings is a natural generalization of the considerations provided herein. Relevant studies are currently in progress and will be reported in future publications.
@@ Acknowledgements.
We are thankful to Victor Ruban for useful discussions and insights. We acknowledge the support from the US National Science Foundation under Grants No. PHY-2110038 (R.C.G.) and No. PHY-2110030 (P.G.K.). W.W. acknowledges supports from the National Science Foundation of China under Grant No. 12004268.

### v2

@@ III.2 Vortex-bright dynamics
We now leverage the ODE reduction dynamics presented in Sec. II.2 to describe the dynamics of the VB in the original PDE model. To generate rotational dynamics of the VB state at the level of the original GP model (1), we displace the stationary solutions away from the origin and seed the waveform into a ground state cloud according to the following procedure. First we compute, using our fixed-point algorithm, the ground state cloud $\psi_{1}=\psi_{\rm gs}$ for a given chemical potential value $\mu_{d}$ and a fixed value of the inter-species coupling $g_{12}$ (while we keep the trap strength constant $\Omega=0.1$ throughout this analysis). For convenience we choose a starting point with $x(0)=x_{0}$ and $y(0)=0$. Next, we evaluate the local chemical potential $\mu_{x_{0}}=|\psi_{\rm gs}(x_{0},0)|^{2}$ where the VB will be seeded. The VB solution with $\mu_{d}=\mu_{x_{0}}$ for a chosen $\mu_{b}$ value is prepared and subsequently imprinted at the location $x_{0}$ onto the ground state cloud $\psi_{\rm gs}$. The resulting displaced structure is propagated according to the dynamical equation (1). Note that to avoid unnecessary oscillations of the background cloud induced by the imprinted VB, the initial condition is “kicked” with a suitable (small) velocity $k$ by boosting $\psi_{2}$ with $e^{iky}$ in order to eliminate any linear momentum for the dark cloud. Our study of the resulting VB trajectories will focus mostly on parameter regimes where flowering trajectories are obtained.
@@
For the relatively small $g_{12}$ values shown in Fig. 1, the resulting PDE trajectories after displacement away from the origin are observed to be approximately circular with additional wobbling motion that does not genuinely resemble flowering trajectories. Nonetheless, for the relatively larger $g_{12}$ values shown in Fig. 2 and for large enough $\mu_{b}$ values, the PDE trajectories result in flowering motion. It is in this regime that we will use the ODE reduction provided in Sec. II to match the full GP VB dynamics. For example, keeping $\Omega=0.1$ and for $g_{12}=1.6$, $\mu_{d}=4.25$, and $x_{0}=2.5$, GP flowering trajectories are seen for $\mu_{b}\gtrsim 4.4$ while spiral-outs are seen for $\mu_{b}<4.4$. This is consonant with the stability results depicted in Fig. 5. However, it is important to note that the stability results of Fig. 5 correspond to a VB that is centered at the origin while we are now seeding VB structures away from the center. For instance, we note that within the $\mu_{b}$-interval $4.44.383<\mu_{b}<$ 4.4 stationary (i.e., centered at the origin) solutions are stable, but the displaced state starting at $x_{0}=2.5$ results in a spiral-out motion. This evidences that this displacement is relatively far away from the origin for the (linear) stability analysis at the origin to be valid. Nonetheless, we have confirmed that, in the unstable region $\mu_{b}\lesssim 4.383$, slightly perturbed solutions from the origin tend to oscillate with frequencies given by Im$(\nu)$.
@@
To compare the ODE reduction with the full GP dynamics in the stable regime of Fig. 5 we also depicted therein the value for $\dot{\theta}$ from the ODE model (see the two cyan curves) overlaid to demonstrate the frequency matching between the PDE and ODE models. This matching is done according to the following procedure. For a given flowering trajectory in the PDE model within appropriate parameter regimes, to obtain the corresponding ODE parameters we first identify those eigenvalues where the Krein signature bifurcates from zero to positive and negative values. The eigenvalues in Fig. 5 are color coded using (blue, black, red) points when the Krein sign is (negative, zero, positive). In particular, for a given $\mu_{b}$ value, we use the upper branch of the bifurcated eigenvalues to obtain the value $\dot{\theta}={\rm Im}(\nu)$. Then, by Eq. (12) we write the corresponding ODE $\Lambda$ parameter as:
$\Lambda=2[\rho(r_{0})\dot{\theta}-M_{0}\dot{\theta}^{2}].$
@@
Figure 9: VB trajectories for $g_{12}=1.6$, $M=0.01$, $\mu_{d}=4.25$, and $\mu_{b}=4.56$. The VB is seeded at $x_{0}=2.5$ corresponding to a local chemical potential $\mu_{x_{0}}\approx 4.22$. To match the ODE trajectories we use $M_{0}\approx 18.9$. The top-left panel shows the resulting trajectory with initial velocities $k_{y}=0$ in the PDE (blue curve) and $\dot{y}(0)\approx 0.01$ in the ODE (red curve) with a resulting angular momentum of $\ell\approx 8.9$. For this value of $\ell$, the effective potential $V_{\rm eff}(r)$ in Eq. (11) is plotted in the bottom-right panel (blue). The three black dots plotted on $V_{\rm eff}(r)$ correspond to its minimum and the minimum and maximum radii of the trajectory. From these values we have an effective flower energy of $\Delta E\approx 0.002$ according to Eq. (16). Parameters are the same in the top-right panel except with initial velocities $k_{y}=0.1$ in the PDE and $\dot{y}(0)\approx 0.06$ in the ODE, with $\ell\approx 8.92$ giving the effective potential in the bottom-right panel (red) with $\Delta E\approx 0.01$. Similarly, the bottom-left panel corresponds to $k_{y}=0.15$ in the PDE and $\dot{y}(0)\approx 0.08$ in the ODE, while $\ell\approx 8.93$ giving the effective potential in the bottom-right panel (yellow) with $\Delta E\approx 0.03$. The thin black circles in the trajectory plots correspond to the circular orbits at the minima of the effective potentials.
@@
Figure 10: Similar to Fig. 9 with the same parameters except here $\mu_{b}\approx 4.47$, $M=0.02$, and $M_{0}\approx 37.0$. The top left panel has initial velocities $k_{y}=0$ in the PDE and $\dot{y}(0)\approx 0.01$ in the ODE, $\ell\approx 8.9$, and $\Delta E\approx 0.006$. For this $\ell$, $V_{\rm eff}(r)$ is in the bottom right panel (blue). For the top right panel $k_{y}=0.1$ in the PDE, $\dot{y}(0)\approx 0.06$ in the ODE, with $\ell\approx 8.95$ to give the effective potential in the bottom right panel (red) with $\Delta E\approx 0.016$. Similarly in the bottom left panel $k_{y}=0.15$ in the PDE, $\dot{y}(0)\approx 0.08$ in the ODE, $\ell\approx 8.97$, and the corresponding effective potential is in the bottom right (yellow) with $\Delta E\approx 0.056$.
@@
Figure 8 depicts the corresponding ODE parameters across a range of $\mu_{b}$ values to the right of the Hamiltonian Hopf bifurcation point. The blue dashed line $P_{2}/\pi$ depicts the estimate for our ODE parameter $M_{0}$ according to Ref. Ruban (2022) [note their $(\tilde{M},M_{br})$ corresponds to our $(M_{0},P_{2})$ respectively as their Eq. (29) agrees with our Eq. (7); see also their Eq. (24) giving $M_{0}=M_{br}/\pi=P_{2}/\pi$]. Notice that our model differs from Refs. Richaud et al. (2021) and Richaud et al. (2022) where they use parameter $\mu$ to represent the mass ratio which is $M=P_{2}/P_{1}$ in our notation. The match of the relevant curves in Fig. 8 is very good, revealing only a slight relative discrepancy of no more than $3.5\%$ for the estimate of $M_{0}$. On the other hand, the red dashed line corresponds to the precessional frequency of a vortex in the absence of the bright component given by $\Lambda=2\omega_{\rm pr}^{0}\mu_{d}$ obtained from Ref. Middelkamp et al. (2010) as follows. Taking our Eq. (7) with $M_{0}=0$ near the origin we obtain
@@
Comparing this reduced system to Eqs. (3) and (4) of Ref. Middelkamp et al. (2010) gives $\Lambda=2w_{\rm pr}^{0}\mu_{d}\equiv\Omega^{2}\ln(8.88\mu_{d}/\Omega)$. The relatively large discrepancy between $\Lambda$ and the theoretical estimate may stem from the fact that the latter, rather than being a true estimate for the precessional frequency of the (massive) VB, it is just the precessional frequency corresponding to a vortex in the absence of the bright component. Note that, for this region of parameters (see Fig. 2), as $\mu_{b}$ increases the mass of the bright component decreases and thus, correspondingly, the $\Lambda$ curve progressively approaches the (red dashed) $\Lambda=2\omega_{pr}^{0}\mu_{d}$ non-filled vortex benchmark. In the two-component VB setting, these results render evident the dependence of $\Lambda$ on the bright component, suitably dressing the effective (parabolic confinement) landscape that the composite VB structure encounters in connection to its motion.
@@ IV Summary and outlook
In the present work we have explored the topic of the dynamics of VB solitary waves deep within the immiscible regime. Our motivation in revisiting this subject stemmed from the renewed experimental ability to monitor such systems which creates the potential for connecting a particle-based understanding with a detailed analysis of the full PDE model. In that vein, we explored equilibrium solutions in this deeply immiscible regime (while some of the earlier studies of the present authors had confined themselves to regimes close to, e.g., hyperfine states of 87Rb where the $g_{ij}$ coefficients are nearly equal Law et al. (2010)). In probing the excitations around the VB equilibrium state, we found the possibility of two frequencies of precession. One of these represents the frequency of precession that is well-familiar from a single-component vortex in a trap. We also identified an additional frequency that can be associated with the oscillation of the VB structure inside an effective radial potential (along the radial direction). The combination of these two frequencies gives rise to the potential for epicyclic motion, which, under suitable conditions discussed herein, can lead to “flowering”, i.e., to flower-like curves of the dynamics of the VB pattern. This intriguing feature, while observed and discussed previously Richaud et al. (2021); Ruban (2022) is here intimately connected between the ODE and PDE dynamics (both at the level of linearization and at that of fully nonlinear dynamical evolution). Our analysis enables both a reasonable approximate analytical description of the relevant motion and key byproducts thereof, such as the critical oscillation amplitude leading to flowering as a (simple) function of the precessional and radial oscillation frequencies. Among the additional intriguing features identified herein, we highlight the (oscillatory) instability of the VB structure when the relevant two frequencies collide. The result of the corresponding dynamics was found to be particularly unusual, in our experience of such systems. Indeed, as the VB embarks on a growing spiral oscillation, it becomes a patch of vorticity that gets elongated and eventually breaks into smaller patches carrying multiple localized phase jumps inside an extended region of vanishing density (in the so-called dark component) in a way reminiscent of persistent currents Wright et al. (2013).
@@
While we considered a number of diagnostics (such as angular momentum-based ones) and visualizations (such as the 3D isocontours of vorticity) of this phenomenon, clearly it merits further study. An additional natural theme of study is to incorporate the (two or more) VB interactions to the model systematically developed herein, so as to examine the potential of formation of lattices (crystals) of such patterns. Understanding the modes of such crystals, their potential instabilities and dynamics promises to be an interesting avenue for the future. It is relevant to also point out here that an alternative formulation of such single and multi-VB configurations is given in Ref. Richaud et al. (2023). A comparative analysis of the latter with the effective particle system produced herein (and its multi-VB generalizations) would be a worthwhile topic for further investigation. Furthermore, the states of interest can be generalized to three spatial dimensions where the natural analogue is the vortex-ring-bright (filled core) solitary wave Ruban et al. (2022); of course, vortex-line-bright structures exist too Charalampidis et al. (2016). A systematic particle-based understanding of such 3D settings is a natural generalization of the considerations provided herein. Relevant studies are currently in progress and will be reported in future publications.
@@ Acknowledgements.
We are thankful to Victor Ruban for useful discussions and insights. We acknowledge the support from the US National Science Foundation under Grants No. PHY-2110038 and No. PHY-2408989 (R.C.G.), No. PHY-2110030, PHY-2408988 and DMS-2204702 (P.G.K.). P.G.K. is also grateful to Profs. Vassilis Koukouloyannis and Ioannis Gkolias for discussions. W.W. acknowledges support from the National Science Foundation of China under Grant No. 12004268.